Two electrons are separated by a distance of 1.60 m. What is the magnitude of the electric force between them?
1
Expert's answer
2020-02-24T10:52:02-0500

Electric charge of electron is:

"q = 1.602 * 10^{-19}C"

"r = 1.6 m"

"k = 9 * 10^9"

The magnitude of the electric force is:

"F = k * \frac{q*q}{r^2}"

"F = 9*10^9*\frac{1.602*10^{-19}*1.602*10^{-19}}{1.6^2}=9.023*10^{-29}N"

Answer: "9.023*10^{-29}N"